RYDBERG STATES CONVERGING TO HIGHER VIBRATIONALLY EXCITED STATES OF HCO^{+}.

\underline{ROBERT J. FOLTYNOWICZ}, WOLFGANG RUPP, HARTMUT G. HEDDERICH and EDWARD R. GRANT, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

Single rotational levels photoselected in the strongest bent B^{2}A′ state of HCO serve as intermediate states in double resonance transitions to Rydberg states built on higher vibrationally excited levels of the linear HCO^{+} core. Frank-Condon factors for the bent-to-linear transitions discriminate against resonances converging to lower lying vibrational states. Widths and intensities of multiple quantum vibrational autoionization resonances are affected by the level of core excitation.